OpenStudy (anonymous):

what is the average kinetic energy of 1 mole of a gas at -32 degrees Celsius?

OpenStudy (anonymous):

The average kinetic energy means the temperature. Use PV=nRT, solve for T. Then you can say Ke=1/2m Vsquared. T=Ke_avg=1/2m Vsquared to expand. Then use E=mc2 and become genius.

OpenStudy (wolfe8):

O.o @Nuge isn't there any easier way to do that?

OpenStudy (anonymous):

Yeah there is. You only need two equations, pV = nRT, where T=Ke_avg. Then set T=Ke=1/2mV^2. Solve for V, you should get some square root, still obtain the same formula. But now you know where this formula comes from. Even easier.